MATH 457: Introduction to Mathematical Logic Homework #5 Due April 2, 2012

We refer to Chapter 4 of the Mathematical Logic lecture notes.

1. Exercise 4.2.5.

2. Exercise 4.2.8 part 1.

3. Exercise 4.2.8 part 2.

4. Exercise 4.2.9.